V=1/3pir2h
the volume is 18.84 and the height is 8 how do i find the radius?

Answer by nerdybill(7384)

V=1/3pir2h
the volume is 18.84 and the height is 8 how do i find the radius?
.
Plug is given into the formula:
18.84 = (1/3)(3.14)r^2(8)
18.84 = (3.14/3)r^2(8)
18.84 = (25.12/3)r^2
18.84(3/25.12) = r^2
56.52/25.12 = r^2
2.25 = r^2
1.5 = r